OpenStudy (anonymous):

Identify the radius of the circle whose equation is (x - 2)2 + (y - 8)2 = 16.

OpenStudy (anonymous):

Plzz help

OpenStudy (welshfella):

compare your equation with the standard one: (x - a)^ + (y - b)^2 = r^2 where r is the radius

OpenStudy (welshfella):

so r^2 = ?

OpenStudy (anonymous):

So the answer is 2?

OpenStudy (welshfella):

sorry that is r^2 = 16 so r = ?

OpenStudy (anonymous):

8

OpenStudy (welshfella):

No r^2 means r * r so you need the square root of 16

OpenStudy (welshfella):

what number multiplied by itself gives 16?

OpenStudy (anonymous):

The answer is 4

OpenStudy (welshfella):

right

OpenStudy (welshfella):

4^2 = 16

OpenStudy (anonymous):

Thank you very much

OpenStudy (welshfella):

the radius of this circle is 4

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

The Eastern Point Lighthouse in Gloucester, Massachusetts was established in 1832 to help fishermen into the harbor. It is still used today by the Coast Guard. A fisherman along the shore wanted to know how tall it was. He is 6 ft. tall and he measures his shadow to be 1.7 ft. The lighthouse casts a shadow of 10.2 ft., approximately how tall is the Eastern Point Lighthouse?

OpenStudy (welshfella):

sorry i've got to go right now

OpenStudy (anonymous):

okay